OpenStudy (anonymous):

What are the exact solutions of x^2 − 5x − 1 = 0?

OpenStudy (abb0t):

Use the quadratic formula for this, since you have: \(ax^2+bx+c\) plug it into: \[\frac{ -b \pm \sqrt{b^2-4ac} }{ 2a }\]

OpenStudy (anonymous):

-5+2^?

OpenStudy (anonymous):

@abb0t

OpenStudy (abb0t):

did you plug in everything correctly? I don't think that looks right.

OpenStudy (anonymous):

is the 5 possitive instead?

OpenStudy (anonymous):

@abb0t

OpenStudy (abb0t):

Yes, it should be but that still looks wrong.

OpenStudy (abb0t):

Did you plug in everything correctly? a=1 b= -5 c = -1

OpenStudy (anonymous):

+5+-2^5/2?

OpenStudy (anonymous):

@abb0t

OpenStudy (abb0t):

\(\huge\frac{ -(-5) \pm \sqrt{5^2-4(1)(-1)} }{ 2(1) }\) does that look like your answer?!?!

OpenStudy (anonymous):

yea

OpenStudy (abb0t):

Lol. no. it doesnt. Lol.